### v1

# Improving Image Captioning Descriptiveness by Ranking and LLM-based Fusion

Abstract
State-of-The-Art (SoTA) image captioning models often rely on the Microsoft COCO (MS-COCO) dataset for training. This dataset contains annotations provided by human annotators, who typically produce captions averaging around ten tokens. However, this constraint presents a challenge in effectively capturing complex scenes and conveying detailed information. Furthermore, captioning models tend to exhibit bias towards the “average” caption, which captures only the more general aspects. What would happen if we were able to automatically generate longer captions, thereby making them more detailed? Would these captions, evaluated by humans, be more or less representative of the image content compared to the original MS-COCO captions? In this paper, we present a novel approach to address previous challenges by showcasing how captions generated from different SoTA models can be effectively fused, resulting in richer captions. Our proposed method leverages existing models from the literature, eliminating the need for additional training. Instead, it utilizes an image-text based metric to rank the captions generated by SoTA models for a given image. Subsequently, the top two captions are fused using a Large Language Model (LLM). Experimental results demonstrate the effectiveness of our approach, as the captions generated by our model exhibit higher consistency with human judgment when evaluated on the MS-COCO test set. By combining the strengths of various SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich, informative nature of human-generated descriptions. This advance opens up new possibilities for generating captions that are more suitable for the training of both vision-language and captioning models [^1] Code and data are available at https://drive.google.com/drive/folders/1tXbJyvwrEF-a-t-NZbt2xjoKaiGGeaz1?usp=sharing..
1 Introduction
Image captioning endeavors to produce organic and human-like descriptions corresponding to a given image. This task holds substantial significance in numerous practical applications, including human-computer interaction and multi-modal recommendation systems. Consequently, it has garnered considerable research interest in recent times. Notably, several State-of-The-Art (SoTA) models have emerged Alayrac et al. (2022); Hu et al. (2022a); Wang et al. (2022b); Yu et al. (2022), demonstrating highly encouraging outcomes when evaluated against widely used metrics such as BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), and SPICE Anderson et al. (2016). These models are typically trained on the Microsoft COCO (MS-COCO) dataset Lin et al. (2014), which contains captions provided by human annotators. MS-COCO has certain limitations: (i) consists of images depicting common objects and scenes, but lacks images of rare or complex events, such as disasters, sports, or art; (ii) captions and queries associated with images often tend to be simplistic, repetitive or inaccurate, which does not reflect the use of natural language and user expectations. Therefore, there is a need for more realistic, large-scale and diverse datasets that can capture the variety and richness of visual and linguistic information in the real-world. The collection of high-quality annotated image-text pair datasets is both time and cost prohibitive. Recently, some vision-language alignment approaches have exploited SoTA captioners for automatically describing the content of images collected from the web Li et al. (2022, 2021a); Wang et al. (2022c). However, the alt-texts often do not accurately describe the visual content of the images, making them a noisy signal that is suboptimal for learning vision-language alignment.
In this paper we propose a captioning model for the creation of high-quality and high-descriptive image-text pair datasets without requiring human intervention. Given an image, our novel training-free model generates captions leveraging several SoTA captioners, then uses a text-image based metric to rank them. The first two captions are finally merged using a Large Language Model (LLM) to generate the final caption. The proposed model will also allow to answer the following Research Questions (RQs):
  • RQ1.
    Is it possible to get more detailed captions by fusing the captions generated by the most advanced SoTA captioning models?
  • RQ2.
    Would the generated captions be more representative of the image content according to human judgement?
  • RQ3.
    Can current image caption models generate richer and more descriptive captions if they are trained using ground-truth data that exhibits these characteristics?
Experimental results demonstrate the effectiveness of our model, as the captions generated by our model exhibit higher consistency with human judgment when evaluated on the MS-COCO test set. By combining the strengths of various SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich, informative nature of human-generated descriptions.
2 Related works
Since the advent of deep learning as the predominant approach, there has been significant advancement in image captioning Anderson et al. (2018); Donahue et al. (2015); Hu et al. (2022a); Huang et al. (2019). Transformer-based architectures have gained particular prominence in recent times Li et al. (2020); Shen et al. (2022). Many research papers follow the strategy of initially pretraining large vision-language models and then adapting them for specific tasks such as captioning Li et al. (2023, 2021b); Wang et al. (2022c). Present efforts mainly revolve around further scaling these pretraining-based methods Alayrac et al. (2022); Hu et al. (2022b); Wang et al. (2022a); Yu et al. (2022) and integrating various vision-language tasks during the pretraining stage Cho et al. (2021); Wang et al. (2022b).
3 Proposed model
Figure 1 shows an overview of the proposed image captioning model. Given $N$ SoTA image captioning models and an image $I$, it is possible to generate $N$ different captions $\{T_{1},...,T_{N}\}$ that describe the content of the image. Generated captions are then ranked according to their alignment to the image content through an Image-Text Matching (ITM) method. Finally, the meaning of the top-2 ranked captions is fused by a LLM to generate a new caption.
Figure 1: Overview of the proposed image captioning pipeline.
3.1 Caption generation
We have carefully selected five cutting-edge image caption model that excel in terms of performance, reproducibility, scalability, and popularity. Our decision to choose five models aligns with the number of ground-truth captions provided by MS-COCO. However, it is important to note that this number can be easily scaled up to accommodate a higher number of models, albeit at an increased computational cost. The selected models include Bootstrapping Language-Image Pre-training 2 (BLIP-2) Li et al. (2023), ExpansionNet v2 (referred to as ExpNet-v2) Hu et al. (2022a), Generative Image-to-text Transformer (GIT) Wang et al. (2022a), One For All (OFA) Wang et al. (2022b), and ViT-GPT2 NLP Connect (2022). For models that have multiple variants, the largest publicly available variant is chosen to ensure greater effectiveness. In particular for BLIP-2 the ViT-g OPT2.7B is used, GITLarge with the ViT-L/14 from CLIP Radford et al. (2021) image encoder is selected, finally OFAHuge exploiting a ResNet-152 as backbone and a Transformer architecture. We point out that the proposed approach is not limited to the number or models that have been chosen for current experiments. Table 1 presents the key characteristics of the selected models. Additional details can be found in the Appendix A.1.
Table 1: Main characteristics of the considered SoTA captioning models.
Model Image encoder Language model/encoder #Param.
BLIP-2 (2023) Li et al. (2023) ViT-g OPT2.7B 3.8B
ExpansionNet v2 (2022) Hu et al. (2022a) SwinT-L FeedForwardNet 236M
GIT (2022) Wang et al. (2022a) ViT-L/14 Transformer 394M
OFA (2022) Wang et al. (2022b) ResNet152+Transformer BPE 930M
ViT-GPT2 (2022) NLP Connect (2022) ViT-B/16 GPT2 302M
In Figure 2, two MS-COCO images are displayed along with the captions generated by the five considered image captioning models. While these captions share a similar length, they vary in terms of their semantic descriptions and level of fidelity. For instance, ExpNet-v2’s caption about the first image states “A herd of sheep standing in a field”, whereas OFA predicts “A herd of sheep with tags in their ears”. Furthermore, in the second image, BLIP-2 describes it as “A group of men standing next to a boat…”, while GIT predicts it as “A group of people in a boat…”.
Image
BLIP-2
ExpNet-v2
GIT
OFA
ViT-GPT2
Refer to caption A herd of black and white sheep in a pen. A herd of sheep standing in a field. A herd of sheep standing on top of a dirt field. A herd of sheep with tags in their ears. A herd of sheep standing in a field.
Refer to caption A group of men standing next to a boat in the water. A group of men standing in the water with a boat. A group of people in a boat on the beach. A group of men standing next to a boat on the beach. People in a boat on the water.
Figure 2: Generated captions for two images of the MS-COCO test set.
3.2 Caption ranking
The goal of this step is to select two captions from the entire set of generated captions $\{T_{1},...,T_{N}\}$. We first define the overall matching score $o_{i}$ of each caption. The captions generated by the chosen models can be evaluated in terms of fidelity, i.e., how much the generated caption is related to the input image without distortion; of adequacy, that is, how much image gist it conveys; fluency, how fluent and beautiful are the language and words Wang et al. (2021). Since the above three aspects are related to both image and text modalities, we use a multi-modal method to estimate the match between each generated caption and the input image. Specifically, we exploit BLIP for Image-Text Retrieval (ITR) Li et al. (2022). The pretrained BLIP is finetuned for ITR by minimizing the Image-Text Matching (ITM) and Image-Text Contrastive (ITC) loss functions. An input image $I$ is encoded by the image encoder into a sequence of embeddings: $\{\mathbf{v}_{\mathrm{cls}},\mathbf{v}_{1},...,\mathbf{v}_{L}\}$, where $\mathbf{v}_{\mathrm{cls}}$ is the embedding of the [CLS] token. The text encoder transforms each generated caption $T_{i}$ into a sequence of embeddings $\{\mathbf{w}_{\mathrm{cls}},\mathbf{w}_{1},...,\mathbf{w}_{M}\}$.
Cosine similarity. The cosine similarity between image-text embeddings is then estimated as $s=g_{v}(\mathbf{v}_{\mathrm{cls}})^{\top}g_{w}(\mathbf{w}_{\mathrm{cls}})$, where $g_{v}$ and $g_{w}$ are linear transformations that map the [CLS] embeddings to normalized lower-dimensional representations.
Matching probability. The image features are fused with the text features through cross attention at each layer of the Image-grounded text encoder. A Fully-Connected (FC) layer followed by softmax predict a two-class probability $\mathbf{p}\in\mathbb{R}^{2}$ indicating the image-text matching probability.
BLIPScore. The overall matching score, referred to as the BLIPScore for brevity, between the image embedding and each generated caption embedding is computed by combining the cosine similarity value and the element of $\mathbf{p}$ representing the matching confidence, $p^{+}$, as $o=\frac{s+p^{+}}{2}$.
Let $\bm{\pi}=\{\pi_{1},\ldots,\pi_{N}\}=\mathrm{argsort}(o_{N})$ be the ranking induced by $o$ for the $N$ generated captions. We then get the two captions, $T_{\pi_{1}}$ and $T_{\pi_{2}}$, which exhibit superior semantic alignment with the provided image, $I$. Figure 3 shows a test image for which the generated captions have been sorted by BLIPScore. As can be seen, the captions generated by BLIP-2 and OFA exhibit the highest BLIPScores, reaching approximately 0.73. They offer complementary descriptions of the image: BLIP-2 focuses on the animals surrounding the book, while OFA detects the presence of a stuffed animal behind the book. Moreover, there is a noticeable difference of approximately 0.50 in the BLIPScores between BLIP-2 and ViT-GPT2. This gap is mainly attributed to an error in object quantification, as BLIP-2 mentions “A bunch of stuffed animals”, whereas ViT-GPT2 identifies only “A stuffed animal…on a couch”.
Model
Caption
Matching probability Cosine similarity BLIPScore
BLIP-2 A bunch of stuffed animals sitting around a book. 0.9907 0.4846 0.7376
OFA A stuffed animal and a book on a chair. 0.9745 0.4891 0.7318
GIT A stuffed animal book with a stuffed animal on it. 0.7402 0.4180 0.5791
ExpNet-v2 Stuffed animals sitting on a couch with a box. 0.5186 0.4272 0.4729
ViT-GPT2 A stuffed animal is sitting on a couch. 0.0689 0.3646 0.2167
Figure 3: Image-text matching scores for the captions generated by using the five considered models on an MS-COCO test image. The captioning models are reported in descending order with respect to the BLIPScore.
3.3 Caption fusion
The first two captions of the rank, namely $T_{\pi_{1}}$ and $T_{\pi_{2}}$, are selected to generate a new caption. The GPT-3 model, also known as Davinci, developed by OpenAI, is utilized for captioning fusion Brown et al. (2020). Davinci stands out as one of the largest and most proficient LLMs currently available. It undergoes training on an extensive corpus of textual data using the unsupervised Transformer learning mechanism, a pre-built learning technique. As a result, the model autonomously learns natural language structures, syntax, rules, and semantics, achieving text generation capabilities comparable to human-like proficiency. In addition to text generation, Davinci demonstrates the ability to analyze sentence syntax and semantics to extract relevant information and generate answers to queries or questions to the best of its abilities.
Hence, to merge the two best generated captions, the following question is prompted to Davinci: “Combine the meaning of these 2 sentences into 1 sentence, considering the semantic meaning and the syntactic meaning. The sentences are: caption1; caption2. These sentences describe an image, I want to get the best caption of the image, using the information in these two sentences.” Here, caption1 and caption2 are $T_{\pi_{1}}$ and $T_{\pi_{2}}$, namely the two captions with the highest BLIPScore. The question to ask Davinci was engineered and the different considerations that lead to the final version can be found in the Appendix A.2.
Figure 4 presents an image from the MS-COCO test set accompanied by the five ground-truth captions, the five captions generated by the SoTA models, and the caption generated by our model. The notable observation is that our generated caption is not only more detailed but also more articulate than competitors. This can be attributed to the advanced generative capabilities of Davinci, which utilizes a larger and more extensive vocabulary compared to conventional image captioning models. The utilization of the Davinci model as a text generator significantly enhances the grammar, syntax, and lexical diversity of the fused caption, thereby producing a remarkably human-like description. For additional qualitative results, please refer to Appendix B.2.
Ground-truth captions:
A metal cup filled with scissors and two ladels.
A silver cup is holding scissors and tongs.
Scissors and other utensils in a cup sitting on a desk.
Metal jar holding scissors, ladles, and tongs on a counter.
A collection of kitchen utensils in a metal bowl.
Captions generated by existing models:
BLIP-2: A metal cup filled with scissors and spoons.
ExpNet-v2: A pair of scissors sitting on top of a metal cup.
GIT: A metal cup holding a pair of scissors and a measuring cup.
OFA: A metal cup filled with scissors and other utensils.
ViT-GPT2: A pair of scissors and a measuring cup on a table.
Caption generated by the proposed model:
Our: A metal cup filled with scissors and other utensils, such as spoons.
Figure 4: Human-written captions versus generated captions. For this sample image of the MS-COCO test set we report the ground-truth captions, those generated by the five existing models, and the one generated by the proposed model.
4 Analysis
Our analysis to verify the effectiveness of the proposed model is performed on the widely used MS-COCO Karpathy split Karpathy and Fei-Fei (2015), which includes 5000 images.
4.1 Quantitative analysis
Quality evaluation.
In order to assess the quality of the generated captions, we employ four widely used image-agnostic evaluation metrics, i.e., BLEU@4 (B@4), METEOR (M), CIDEr (C), and SPICE (S). Additionally, we exploit the defined BLIPScore to further evaluate the performance of the generated captions. The BLIPScore allows to estimate the semantic alignment between the generated captions and the provided image.
Diversity evaluation.
To investigate the diversity of the generated captions, we use mBLEU, and n-gram diversity (i.e., Div-$n$ Aneja et al. (2019)). These metrics evaluate diversity by comparing the $n$-gram differences among the generated captions that belong to the same image.
Richness evaluation.
Part-of-Speech (POS) plays a crucial role in analyzing text semantics. To compare the semantic richness of the captions generated by the different models, we adopted POS tagging. In this regard, the Universal tagset consisting of 17 English POS tags is exploited Smith and Eisner (2005). Initially, the words in each caption are annotated with their respective POS tags. Subsequently, we estimate the frequency of each POS tag within the captions generated by each model.
Results.
In our analysis we include the five literature models considered for caption generation, namely BLIP-2, ExpNet-v2, GIT, OFA, and ViT-GPT2. The performance of these models is compared with that of our proposed model, denoted as “Our (Fusion)”. Furthermore, we evaluate the performance achieved by selecting the caption with the highest BLIPScore from the five generated captions during the ranking phase for each image. The results obtained using this approach are presented as “Our (Best)”.
Several observations can be drawn from the quality and diversity results presented in Table 2. First, the performance of our proposed model in terms of captioning quality metrics is either on par with or slightly lower than existing models. It is worth noting that there is a noticeable decline in performance compared to the best-performing method, which is OFA, specifically for metrics B@4 and C, where our model lags behind with scores of 14 and 50, respectively. However, our model achieves comparable results to OFA for metrics M and S. Second, the proposed model is able to generate captions that describe the image more adequately. The BLIPScore is in fact 0.05 higher than the first SoTA model, which is OFA. We also highlight that the “Our (Best)” approach also obtains a BLIPScore equal to that of “Our (Fusion)”. In Appendix B.1, we provide a detailed analysis of the BLIPScore, including statistics on the models to which the selected captions for the fusion step belong. Third, diversity metrics demonstrate how the proposed model generates captions that are very different from the others. This difference is particularly noticeable in the case of Div-1 and Div-2, where the gap with the leading SoTA captioner is 0.15. Regarding mBLEU, our proposed model ranks second to ViT-GPT2. However, it is worth noting that ViT-GPT2 achieves higher diversity at the expense of faithfully and fluently describing the content, as evidenced by its significantly lower quality captioning performance.
Table 2: Captioning performance in terms of image-agnostic metrics together with diversity statistics on the Karpathy split of MS-COCO. For each metric, best and second best models are highlighted in bold and underlined, respectively.
Model Quality Diversity
B@4 ($\uparrow$) M ($\uparrow$) C ($\uparrow$) S ($\uparrow$) BLIPScore ($\uparrow$) mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$)
BLIP-2 43.8 31.7 145.9 25.2 0.64 0.88 0.50 0.32
ExpNet-v2 40.7 30.0 139.5 24.4 0.59 0.84 0.49 0.32
GIT 38.7 29.5 131.2 23.3 0.62 0.80 0.49 0.32
OFA 44.6 32.5 153.7 26.6 0.65 0.91 0.51 0.33
ViT-GPT2 35.4 27.9 119.1 21.2 0.48 0.69 0.47 0.30
Our (Best) 41.6 31.6 145.5 25.9 0.70 1.00 0.52 0.34
Our (Fusion) 30.0 31.5 106.0 26.5 0.70 0.72 0.66 0.44
Table 3 shows the results of the comparison between the captioning models in terms of average number of tokens and frequency of POS-Tags. The first consideration that can be made is that the captions generated by our model are on average longer both than those generated by SoTA models and those composed by humans. Although this finding alone may not hold significant meaning, a closer examination of POS-Tag frequencies, particularly adjectives, nouns, and verbs, reveals that our captions exhibit a greater richness and descriptive quality.
Table 3: Number of tokens and frequency of POS tags for generated captions. Each column reports each captioner’s statistics for the 5000 images of the MS-COCO test set (tags that never occurred are omitted). The “Ground-Truth” column indicates statistics for human generated captions. For each tag, best and second best models are highlighted in bold and underlined, respectively.
BLIP-2 ExpNet-v2 GIT OFA ViT-GPT2 Our (Best) Our (Fusion) Ground-Truth
Token 9.62$\pm$1.42 9.57$\pm$1.75 9.60$\pm$1.59 9.81$\pm$1.28 9.17$\pm$1.53 10.05$\pm$1.57 13.10$\pm$3.32 10.43$\pm$2.36
Adjective 0.62$\pm$0.81 0.52$\pm$1.12 0.51$\pm$0.74 0.55$\pm$0.75 0.39$\pm$0.68 0.66$\pm$0.81 0.91$\pm$0.98 0.85$\pm$0.91
Adposition 1.67$\pm$0.85 1.74$\pm$0.83 1.67$\pm$0.85 1.71$\pm$0.84 1.54$\pm$0.79 1.74$\pm$0.86 2.10$\pm$1.11 1.68$\pm$0.95
Adverb 0.01$\pm$0.10 0.00$\pm$0.03 0.01$\pm$0.10 0.00$\pm$0.07 0.01$\pm$0.07 0.01$\pm$0.09 0.06$\pm$0.26 0.07$\pm$0.28
Conjunction 0.20$\pm$0.41 0.21$\pm$0.44 0.23$\pm$0.44 0.24$\pm$0.45 0.23$\pm$0.43 0.27$\pm$0.47 0.46$\pm$0.59 0.24$\pm$0.47
Determiner 2.37$\pm$0.70 2.50$\pm$0.70 2.41$\pm$0.72 2.45$\pm$0.76 2.40$\pm$0.72 2.44$\pm$0.77 2.93$\pm$1.02 2.21$\pm$0.95
Noun 3.51$\pm$0.85 3.59$\pm$0.84 3.52$\pm$0.90 3.64$\pm$0.81 3.43$\pm$0.82 3.71$\pm$0.89 4.59$\pm$1.42 3.69$\pm$1.15
Numeral 0.09$\pm$0.29 0.11$\pm$0.32 0.09$\pm$0.29 0.14$\pm$0.35 0.05$\pm$0.22 0.14$\pm$0.35 0.16$\pm$0.42 0.12$\pm$0.35
Particles 0.18$\pm$0.40 0.13$\pm$0.35 0.15$\pm$0.39 0.15$\pm$0.36 0.13$\pm$0.36 0.16$\pm$0.39 0.19$\pm$0.42 0.19$\pm$0.43
Pronouns 0.09$\pm$0.29 0.06$\pm$0.24 0.12$\pm$0.35 0.09$\pm$0.28 0.09$\pm$0.29 0.09$\pm$0.31 0.14$\pm$0.37 0.14$\pm$0.38
Verb 0.85$\pm$0.61 0.71$\pm$0.57 0.89$\pm$0.72 0.85$\pm$0.67 0.90$\pm$0.75 0.84$\pm$0.68 1.55$\pm$0.95 1.21$\pm$0.89
4.2 Subjective study
We conduct a subjective study to collect human judgments for selecting the best caption. Specifically, participants were instructed to choose the caption that accurately describes the image, is grammatically correct, contains no incorrect information, is relevant to the image, and is human-like. Participants could select from six different captions generated by using BLIP-2, ExpNet-v2, GIT, OFA, ViT-GPT2, and the proposed model. To avoid bias, captions are randomly arranged each time the page is loaded. We use the Amazon Mechanical Turk (AMT) to gather the human judgements. From the MS-COCO test set, a subset of 1000 images is randomly sampled for evaluation. We conduct two distinct studies: the first involves rating each image by three unique workers, who are not necessarily domain experts; in the second study, we enlist the expertise of three domain experts to rate the images. Figure 4(a) displays a screenshot of the AMT interface designed for caption assessment. The interface is self-contained as instructions to guide the selection are given in the evaluation of each image. Table 4(b) shows the percentage of votes obtained by each model from generic workers (shown in the “Worker votes” column), expert workers (reported in the column “Expert votes”), and the average percentage between the two categories of participants (see column “Average votes”). Our captions have been selected more frequently by both categories of participants with a percentage of 37.77. The second best is OFA with a percentage of 16.18. Non-expert workers show a uniform distribution in the selection of SoTA captions (on average about 15%), while instead they show a greater, albeit limited, preference for our captions with a percentage of 24.33. It is worth noting that experts have consistently chosen our captions, with 51.20% of the votes against an average of about 10%.
Model Worker votes (%) Expert votes (%) Average votes (%)
BLIP-2 15.23 12.97 14.10
ExpNet-2 14.80 6.47 10.63
GIT 16.41 9.93 13.17
OFA 16.73 15.63 16.18
ViT-GPT2 12.50 3.80 8.15
Our 24.33 51.20 37.77
Figure 5: Human judgments interface and model votes distribution. (a) showcases the user interface utilized for collecting human judgments, providing an intuitive platform for gathering evaluations. (b) presents the distribution of votes obtained by each model, illustrating the collective assessments rendered by the participants.
5 Evaluation
In this section, we investigate whether the best SoTA model, specifically OFA, can learn to generate more descriptive captions when trained using our generated captions as groun-truths. To conduct this experiment, we finetune the model for the image captioning task on MS-COCO by replacing the original ground-truths with the captions generated by our model. Due to limited computational resources (see Appendix A.3 for details), we utilize the Base version of OFA and follow the setup parameters provided by the authors. The 5000 samples from the MS-COCO test set are randomly divided into training (3500 samples), validation (500 samples), and testing (1000 samples). To minimize bias introduced by partitioning, we repeat this split process five times. Table 4 reports the quality captioning performance for OFABase trained on the original MS-COCO and finetuned on MS-COCO with our ground-truths, respectively. It is possible to see that the results for the finetuned model are slightly higher than the baseline ones, i.e. 2 for B@4, 1 for M and 10 for C, respectively.
Table 4: Captioning performance comparison between two versions of OFABase, i.e. the model pretrained on MS-COCO (as known as the baseline) and the model finetuned on the dataset with our generated captions used as ground-truth. It is important to note that the reported performances represent the mean and standard deviation across 5 train-val-test splits.
Model B@4 ($\uparrow$) M ($\uparrow$) C ($\uparrow$) S ($\uparrow$) BLIPScore ($\uparrow$)
Baseline 29.06$\pm$0.83 28.75$\pm$0.46 268.96$\pm$9.06 46.88$\pm$0.58 0.6367$\pm$0.0030
Finetuned 31.22$\pm$0.91 29.88$\pm$0.45 281.62$\pm$10.43 46.72$\pm$0.91 0.6509$\pm$0.0018
6 Discussion
Existing datasets and caption models have several limitations, such as short ground-truth captions and generated captions that tend to look alike and are semantically poor. In this article we propose a caption model that tries to overcome the previous limitations by generating longer and more detailed captions without requiring human intervention. The captions generated by our model were evaluated from different points of view. Here, we summarize the contributions of our work, answering the research questions posed in the introduction. In section 4.1, we address [RQ1] and demonstrate that the captions generated by our proposed model outperform those generated by the SoTA models in terms of accuracy, diversity and semantic richness. In section 4.2, we tackle [RQ2] through a subjective study involving human participants. Based on human judgment, our model generates captions that are more appealing than those produced by SoTA models. In section 5, we aim to answer [RQ3]. The results of our experiment highlight how current image captioning models lean towards the use of common words, phrases and linguistic patterns. Consequently, this leads to a dearth diversity in the generated captions, even when the models are trained on more elaborate and descriptive caption datasets Chen et al. (2022). This phenomenon, known as the mode collapse problem, has been widely discussed in generative modeling Zhou et al. (2020). Addressing this issue requires further investment. Consequently, our proposed caption model can serve as a valuable resource for generating ground-truth captions and pushing models beyond their current limitations.
Societal impact.
Compared to previous image captioning models, our model demonstrates improvements in performance and is more suitable for helping people with visual impairments. The models that are exploited by our proposal are pretrained on large-scale data, and the data are not guaranteed to contain no toxic contents, which could negatively impact the output generated. While we have observed only a small number of such cases in our qualitative analysis, it is imperative to exercise caution when deploying the model into practical applications. Furthermore, more research investigations are needed to ensure better control over the model output.
Limitations.
The limitations of the proposed model are mainly three. The first is that sometimes the five SoTA captioners generate identical or very similar captions (this occurs for about 150 of the 5000 images in the MS-COCO test set). In this case the fusion of the captions collapses into a caption that does not offer more detail. An example is the following where BLIP-2, ExpNet-v2 and OFA generate the same caption, namely “a giraffe standing in the middle of a field”, GIT describes the image content as “a giraffe standing in a field with trees in the background”, finally ViT-GPT2 predicts that “a giraffe standing in a field with trees”. The caption generated by our model is “a giraffe stands majestically in the middle of a field”. Secondly, for 16 test images, the LLM model generates captions with the prefix “The caption for the image could be:”. It is important to note that this structural pattern does not result from the fusion of the two captions. Rather, it is a deliberate prompt used by the Davinci model to provide context and generate relevant responses. The third aspect is not a true limitation of the proposed model but rather one of its unique characteristics. Some of the captions generated by the LLM model incorporate emotional elements, adding depth and richness to the captions. For example, phrases like “a trolley car driving down a city street at night creates a captivating scene” or “a green bench in a subway station with a yellow train in the background creates a vivid and vibrant scene”. These emotional aspects are not explicitly provided in the input captions but are instead generated by the LLM model.
7 Conclusion and Future work
This paper presents a training-free image captioning model that can generate highly descriptive captions. By leveraging the information from the best captions produced by existing models, our proposal has demonstrated its ability to generate captions that surpass the accuracy, semantic richness, and appeal of those generated by current models. These findings obtained on the MS-COCO dataset suggest the potential of utilizing our proposed model to generate ground-truth captions for the development of vision-language and captioning models.
In light of the limitations observed in the proposed model, a future direction of development would involve addressing the issue of mode collapse. This could be achieved by enhancing the caption ranking policy, moving beyond solely selecting captions that closely align with the image, and instead prioritizing captions that are diverse from one another. By promoting diversity among the selected captions, we can mitigate the problem of mode collapse and further enhance the richness and variety of our captions.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We describe the contributions of the presented work in the abstract and at the end of Section 1.
    2. (b)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Yes and we believe that our paper conforms to them.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] .
    4. (d)
      Did you describe the limitations of your work? [Yes] We dedicate a paragraph in Section 6 describing the limitations of our proposal.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] All the necessary code is available at this link https://drive.google.com/drive/folders/1tXbJyvwrEF-a-t-NZbt2xjoKaiGGeaz1?usp=sharing
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Yes. For the experimental results of finetuning OFA across multiple data splits we reported the standard deviation.
    4. (d)
      Did you include the amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A.3.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite the articles in which the exploited methods were presented and in Appendix A.1 we report links and details of the methods.
    2. (b)
      Did you mention the license of the assets? [Yes] In Appendix A.1 we mention the license of the used assets.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section 6.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] We describe the purpose of the subjective study and provide a screenshot of the used interface in Section 4.2.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Cost and time estimation for subjective studies are discussed in Appendix B.4.
Appendix A Implementation Details
In this section we describe the selected models used for caption generation. Next we discuss the choices that led to the definition of the prompt for Davinci. Finally, the resources used for the development of the proposed work are detailed.
A.1 Captioning models
BLIP-2 Li et al. (2023) is a vision-language model that leverages pretrained Visual Language Processing (VLP) and NLP models to effectively handle tasks. It employs a Querying Transformer (Q-Former) to connect the image encoder and Large Language Model (LLM) and extract relevant features based on queries. The model undergoes two pretraining phases: Vision-and-Language Representation Learning and Vision-to-Language Generative Learning. The former aligns image and text representations through contrastive learning, grounded text generation, and matching, while the latter connects the Q-Former and frozen LLM. BLIP-2 achieved its best performance using ViT-g/14 as the Image Encoder and OPT2.7B as the LLM for image captioning on the MS-COCO dataset. ExpansionNet v2 Hu et al. (2022a) is an image captioning model that leverages expansion methods for improved performance. The model architecture is based on an encoder-decoder architecture built on top of a Swin Transformer in the Large configuration (SwinT-L) Liu et al. (2021). In this model, the image undergoes encoding using SwinT, followed by Block Static Expansion. During training, the ground-truth caption is encoded using a FeedForwardNet, which incorporates a Dynamic Expansion module. Visual and text representations are then combined using Multi-Head Attention and further processed by a classifier. The model parameters are efficiently learned using an end-to-end alternating full and partial training strategy. GIT Wang et al. (2022a) is a vision-language model that combines an image encoder and a text decoder. The image encoder is based on a Swin-like vision transformer pretrained on image-to-text data. It produces normalized and projected features, serving as input to the text decoder. The text decoder, a Transformer network, generates text descriptions using self-attention and feed-forward layers. The input text is processed and combined with visual features for decoding. Pretraining involves mapping images to their associated text descriptions, using a language modeling objective. The model was trained on 800M image-text pairs, employing a language modeling task for both pretraining and finetuning. OFA Wang et al. (2022b) is a versatile model designed for various vision-language tasks. It follows a task- and modality-agnostic approach, supporting both unimodal and multi-modal tasks. Image and text representations are unified in a token-based space, including object tracking and coordinates. Specifically, an image is encoded with ResNet modules to obtain patch-level embeddings, a text sequence is first tokenized using Byte-Pair Encoding (BPE) Sennrich et al. (2016) then subword tokens are encoded to embeddings. The encoder-decoder architecture consisting of Transformers is pretrained on vision-language, vision, and language datasets using cross-entropy loss. For image captioning, two finetuning steps optimize cross-entropy loss and CIDEr metric. ViT-GPT2 NLP Connect (2022) is an image captioner consisting of a ViT as the image encoder and the GPT2 as the text decoder. ViT processes images by dividing them into patches using attention mechanisms, while GPT2 analyzes and generates natural language text. Both models are pretrained, making implementation straightforward and achieving good performance. The ViT-GPT2 model combines the strengths of these unimodal models, providing a simple yet effective multi-modal approach for image captioning tasks.
The source code for BLIP-$2$https://github.com/salesforce/LAVIS (last access: 14 June 2023) is released under the BSD-3-Clause license, the source codes for ExpansionNet-v2333https://github.com/jchenghu/ExpansionNet_v2.git (last access: 14 June 2023), OFA [^4] https://github.com/OFA-Sys/OFA.git (last access: 14 June 2023), GIT [^5] https://github.com/huggingface/transformers.git (last access: 14 June 2023), and ViT-GPT2666https://huggingface.co/nlpconnect/vit-gpt2-image-captioning/tree/main (last access: 14 June 2023) are available under the Apache-2.0 license. The MS-COCO dataset is distributed under a CC-BY 4.0 license.
A.2 The Davinci model
The Davinci model is integrated into our captioning model using the OpenAI API [^7] https://openai.com/blog/openai-api. Interaction with the API is done through HTTP requests via official Python bindings. The OpenAI API requires API keys for authentication [^8] https://platform.openai.com/account/api-keys.
We conduct several tests to determine the optimal prompt for the Davinci model. It was discovered that even a slight change in certain terminologies can lead to variations in the obtained results, particularly in the case of descriptive captions. Based on the performed tests, the following conclusions can be drawn:
  • Including the information that caption1 and caption2 pertain to the same input image enhances the analysis and aids the model in recognizing shared elements. For instance, if one caption mentions the subject “child” while another refers to the subject as a “young girl”, the model, understanding that both terms denote the same entity depicted in the image, will retain the more precise description, namely “young girl”.
  • Explicitly instructing the model to fuse sentences both syntactically and semantically facilitates the creation of a result that goes beyond a mere summary of the two captions. Instead, it produces a genuine amalgamation of information while eliminating redundancy in terms and/or content. For instance, if both analyzed sentences feature the term “kite” in a specific syntactic role, the model comprehends that there is a single subject and avoids duplicating the term “kite” when generating the final caption.
  • Incorporating the terms “information” and “meaning” within the prompt significantly enhances the quality of the generated captions. They transcend being a mere compilation of terms found in the two analyzed captions, instead resulting in a truly integrated and cohesive generated caption.
A.3 Resource used
To perform the experiments for this work, namely running SoTA captioners and finetuning the OFA model, we exploit a desktop computer with an Intel Core i7-7700 CPU@3.60GHz, 16 GB DDR4 RAM 2400 MHz, and NVIDIA Titan X Pascal with 3840 CUDA cores. The OpenAI API is exploited for the caption fusion step of the proposed model.
Appendix B Analysis
In this section, we begin by conducting an analysis of BLIPScore, followed by presenting qualitative results to highlight the potential of the proposed captioning model. As part of an ablation study, we evaluate the performance of the proposed model using the base version, wherever available for SoTA models. Lastly, we delve into the configuration and outcomes of the study conducted with AMT.
B.1 BLIPScore analysis
In this analysis, we examine the distribution of BLIPScores across the 5000 test images from MS-COCO. Figure 6 displays several boxes representing different captioning models. These include five SoTA models, captions generated by our proposed model, captions with the highest BLIPScore among the five SoTA captioners, and human-made captions. The plot reveals several noteworthy findings. First, our proposed model achieves a slightly higher median score compared to the “Our (Best)” solution, indicating that our model incorporates additional details that enhance the alignment between the description and the image. Second, the ViT-GPT2 model performs the poorest, displaying the least accurate representation of image content. Lastly, it is interesting to observe the high interquartile range of BLIPScores for ground-truth captions. This characteristic arises due to the substantial variability in the estimated BLIPScores. Further examination of images with low BLIPScores revealed that some user-written captions lack detail and are inadequate.
Figure 6: Boxplot presenting the BLIPScore distribution for the 5000 test images with respect to generated and ground-truth captions.
In Figure 7, we present the percentage of SoTA model pairs selected during the ranking phase. In practice, these results demonstrate which models most frequently belong to the captions that are fused by our model. As shown, the pair of models BLIP-2 and OFA is selected for 22% of test image captions, while GIT and OFA are selected for 19%. The pair consisting of ExpNet-v2 and ViT-GPT2 accounts for only 2.4% of the generated captions.
Figure 7: Percentage of MS-COCO images for which caption model pairs were selected based on BLIPScore. For the sake of brevity, only the initials of the model names are mentioned. “B” is for BLIP-2, “E” is ExpNet-v2, “G” is GIT, “O” stands for OFA, and “V” is ViT-GPT2.
B.2 Qualitative results
Figure 8 showcases images accompanied by their respective ground-truth and generated captions. From the provided examples, it is evident that our model excels in producing more detailed captions. For instance, in the image featuring a child with a toothbrush, our model seamlessly integrates information about the toothbrush, overalls, and hands. Similarly, in the image of the tower with a clock, our model enriches the description with intriguing qualitative details such as the presence of red and gold. On the other hand, as consistent with the discussions presented in the section outlining the limitations of our model, when the captions generated by the SoTA models exhibit significant similarity, our generated captions may not provide additional details (as observed in the image featuring the pole on the sidewalk). In some cases, the generated caption might also include redundant information, as seen in the description of the ducks in the pond, where terms like “muddy pond” and “brown water” are used together.
Refer to caption Ground-truth captions: A street sign for queen street with a colorful representation of the street on the pole. A decorated street sign for Queen Street with two blue signs below. A close up a street pole with a homemade street sign. We are looking at a street sign at the corner. A pole holding the street sign for Queen Street is decorated with a painting of a queen. Captions generated by existing models: BLIP-2: A street sign on a pole on the side of the road. ExpNet-v2: A wooden pole with street signs on the side of it. GIT: A street sign on a pole with a blue sign on it. OFA: A street sign on a pole on the side of a street. ViT-GPT2: A street sign on a pole on a sidewalk. Caption generated by the proposed model: Our: A wooden pole with street signs mounted on the sidewalk.
Refer to caption Ground-truth captions: A infant holding a baby toothbrush in his hand looking at it. A small baby is holding a white and blue toothbrush. A child holds a toothbrush in their hand. The baby is holding and looking at his tooth brush. A baby holds a toothbrush in its hand. Captions generated by existing models: BLIP-2: A baby in overalls holding a toothbrush. ExpNet-v2: A baby holding a toothbrush in his hand. GIT: A baby holding a toothbrush in his hands. OFA: A baby holding a toothbrush in its hands. ViT-GPT2: A baby holding a toothbrush in its mouth. Caption generated by the proposed model: Our: A baby in overalls is holding a toothbrush in his hands.
Refer to caption Ground-truth captions: Two Mallards float and drink water in a shallow pool. Two ducks in a pond in a camp grounds. Two ducks swimming in a body of water. Two ducks swimming on a very dirty pond. Two ducks are swimming in some brown water. Captions generated by existing models: BLIP-2: Two ducks swimming in a pool of brown water. ExpNet-v2: Two ducks swimming in the water in a pond. GIT: Two ducks swimming in a muddy pond with a car in the background. OFA: A couple of ducks swimming in a body of water. ViT-GPT2: A duck swimming in a pond with a duck in it. Caption generated by the proposed model: Our: Two ducks swimming in a muddy pond with a car in the background, amidst a pool of brown water.
Refer to caption Ground-truth captions: A large tall tower with a clock on the top. A red, white and gold decorative pillar with a clock on top. Pedestrians rest at the base of a large, public ornamental clock. A large tower in a crowded area with a clock at the top. A tall red and white clock tower in the middle of a street. Captions generated by existing models: BLIP-2: A red and gold tower with a clock on top. ExpNet-v2: A red clock tower in the middle of a street. GIT: A large clock tower with a large clock on it. OFA: A red clock tower with people sitting around it. ViT-GPT2: A clock tower with a clock on it’s side. Caption generated by the proposed model: Our: A red and gold clock tower with people sitting around it.
Figure 8: Human-written captions versus generated captions. For these images of the MS-COCO test set we report the ground-truth captions, those generated by the five existing models, and the one generated by the proposed model.
B.3 Evaluation of base model versions
The captioning pipeline designed was also implemented using the reduced versions of the SoTA image captioning models, i.e. the Base versions of GIT and OFA. As done for the more expensive model versions, also in this case the evaluation was conducted on the MS-COCO test set. Table 5 contains the quality and diversity results of the generated captions. As expected, the performance achieved using the reduced versions of the SoTA models is lower compared to the results reported in Table 2. This difference is particularly noticeable in the captioning metrics. However, it is worth noting that the BLIPScore remains almost identical, with a value of 0.65 for OFAHuge and 0.62 for OFABase. The BLIPScore obtained by our model with the best SoTA versions is 0.70 and 0.69 with the Base versions. These findings suggest that in terms of caption adequacy, the variance between the two model versions has minimal impact.
Table 5: Captioning performance by involving the Base version of GIT and OFA. Performance is reported in terms of image-agnostic metrics together with diversity statistics on the Karpathy split of MS-COCO. For each metric, best and second best methods are highlighted in bold and underlined, respectively.
Model Quality Diversity
B@4 ($\uparrow$) M ($\uparrow$) C ($\uparrow$) S ($\uparrow$) BLIPScore ($\uparrow$) mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$)
BLIP 39.8 31.0 133.4 23.8 0.62 0.78 0.47 0.30
ExpNet-v2 40.7 30.0 139.5 24.4 0.59 0.73 0.44 0.28
GITBase 36.7 28.5 122.9 22.0 0.59 0.70 0.43 0.28
OFABase 42.6 31.3 146.4 25.5 0.62 0.80 0.46 0.30
ViT-GPT2 35.4 27.9 119.1 21.2 0.48 0.59 0.42 0.27
Our (Best) 38.3 30.6 134.3 24.6 0.68 1.00 0.47 0.30
Our (Fusion) 28.0 30.9 98.5 25.6 0.69 0.50 0.60 0.40
B.4 AMT analysis
We had no experience with AMT prior to this study. Initially, we enlisted generic workers to rate each of the 1000 images, compensating them with \$0.02 per evaluation. The remuneration was determined based on the assumption that each assessment would require a maximum of 90 seconds. In total, we allocated \$90 for collecting assessments from the generic workers. We later acquired the judgments from experienced workers who we paid \$0.10 for a total of \$330.
Figure 9 illustrates the level of agreement among the six workers involved in the subjective study when selecting the best caption for each image. The graph addresses the question of how many of the 1000 images received a unanimous consensus among the workers. As observed from the results, there is a general consensus among the workers that our captions are superior. Specifically, for 308 images, two workers agreed on our caption as the best, for 266 images, three workers agreed, for 118 images, four workers agreed, finally, for 31 images, five workers chose our caption.
Figure 9: Level of agreement among the six workers involved in the subjective study.